find the largest 6 digit number divisible by 18,24,36

Grade:10

1 Answers

Arun
6 years ago
First find the LCM of 18 , 24 and 36
18=2×9
=2×3×3
24=2×12
=2×2×6
=2×2×2×3
36=2×18
=2×2×9
=2×2×3×3
LCM=2×2×2×3×3×3
=8×27
=216
We need 6 digit greatest number so muliply it my a number
216×4629=999864